# Towards Unifying Interpretability and Control: Evaluation via Intervention

### v1

@@ Abstract
With the growing complexity and capability of large language models (LLMs), a need to understand model reasoning has emerged, often motivated by an underlying goal of controlling and aligning models. While numerous interpretability and steering methods have been proposed as solutions, they are typically designed either for understanding or for control, seldom addressing both, with the connection between interpretation and control more broadly remaining tenuous. Additionally, the lack of standardized applications, motivations, and evaluation metrics makes it difficult to assess these methods’ practical utility and efficacy. To address the aforementioned issues, we propose intervention as a fundamental goal of interpretability and introduce success criteria to evaluate how well methods are able to control model behavior through interventions. We unify and extend four popular interpretability methods—sparse autoencoders, logit lens, tuned lens, and probing—into an abstract encoder-decoder framework. This framework maps intermediate latent representations to human-interpretable feature spaces, enabling interventions on these interpretable features, which can then be mapped back to latent representations to control model outputs. We introduce two new evaluation metrics: intervention success rate and the coherence-intervention tradeoff, designed to measure the accuracy of explanations and their utility in controlling model behavior. Our findings reveal that (1) although current methods allow for intervention, they are inconsistent across various models and features, (2) lens-based methods outperform others in achieving simple, concrete interventions, and (3) interventions often compromise model performance and coherence, underperforming simpler alternatives, such as prompting, for steering model behavior and highlighting a critical shortcoming of current interpretability approaches in real-world applications requiring control. Code is made available for replicability.
@@ 1 Introduction
As large language models (LLMs) have become more capable and complex, there has emerged a need to better understand and control these models to ensure their outputs are safe and human-aligned. Many interpretability methods aim to address this problem by analyzing model representations, attempting to understand their underlying computational and reasoning processes in order to ultimately control model behaviour. While many of these methods, and interpretability as a field more broadly, claim control and intervention as abstract goals and present compelling qualitative results demonstrating that intervention is possible (for example, Anthropic’s Golden Gate Claude Anthropic ; Templeton et al. (2024)), the link between interpretation and intervention is often tenuous in practice, as many methods are not explicitly tailored for both. Furthermore, even fewer are thoroughly and systematically evaluated for the ability to control model outputs beyond simple qualitative examples. We believe the reason for this is threefold. First, interpretability methods produce explanations in disparate feature spaces, such as the token vocabulary, probe predictions, or learned auto-interpreted features, hindering comparisons across methods. Second, there exists a “predict/control discrepancy” (Wattenberg & Viégas, 2024), where the features identified by interpretability methods for predicting behavior are not the same as those used for steering it. Third, there do not exist standard systematic benchmarks to measure intervention success.
@@
In this work, we view intervention as a fundamental goal of interpretability, and propose to measure the correctness of interpretability methods by their ability to successfully edit model behaviour. In particular, we focus on sparse autoencoders  (Cunningham et al., 2023; Rajamanoharan et al., 2024a; Templeton et al., 2024; Bricken et al., 2023; Gao et al., 2024), Logit Lens  (nostalgebraist, ; Dar et al., 2023), Tuned Lens  (Belrose et al., 2023; Din et al., 2023), and linear probing  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022), and benchmark them with steering vectors and prompting as baselines for intervention. In order to enable comparison across these various methods, we first unify and extend the methods as instances of an abstract encoder-decoder framework, where each method encodes uninterpretable latent representations of language models into human-interpretable features and the decoder of the framework inverts this mapping, allowing us to reconstruct a latent representation from the features. Under this abstract framework, we can intervene on the interpretable feature activations generated by each method and decode them into latent counterfactuals, which produce counterfactual outputs corresponding to the desired intervention.
@@
The unifying feature interpretation and intervention framework allows us to propose two standard metrics for evaluating mechanistic interpretability methods: (1) intervention success rate, which measures how well intervening on an interpretable feature causally results in the desired behavior in the model outputs, and (2) coherence-intervention tradeoff, which measures how well the causal interventions succeed without damaging the coherence of the model’s outputs. We evaluate Logit Lens, Tuned Lens, sparse autoencoders, and linear probes, for these metrics on GPT2-small, Gemma2-2b, and Llama2-7b, comparing them to simpler but uninterpretable baselines of steering vectors and prompting. Our results show that while existing methods allow for intervention, they are inconsistent across features and models. Furthermore, lens-based methods outperform all other methods including sparse autoencoders for simple, concrete features. This is likely due to frequent mislabeling of the learned SAE features, where a feature labeled “references to coffee” actually seems to encode for ‘beans’ and ‘coffee beans’ more than ‘coffee’. Similarly, probes and steering vectors are prone to learning spurious correlations in their training data, whereas Logit Lens and Tuned lens do not suffer from the same sensitivities. Finally, we show that intervention often comes at the cost of model output coherence, underperforming simple prompting baselines, presenting a critical shortcoming of existing methods in real-world applications that require control and intervention. We conclude this work with detailed takeaways about the strengths and weaknesses of each method, including discussion of which methods are optimal for specific intervention topics, which are best to use out of the box, and which hold the most promise for future development.
@@
  • In Section 3.1, we present an encoder-decoder framework that unifies four popular mechanistic interpretability methods: sparse autoencoders, logit lens, tuned lens, and probing.
  • In Section 3.2, we propose two evaluation metrics for encoder-decoder interpretability methods, namely (1) intervention success rate; and (2) the coherence-intervention tradeoff to evaluate the ability of interpretability methods to control model behavior. We also provide a dataset for benchmarking state-of-the-art LLMs on these metrics.
  • In Section 4, we perform experimental analysis on GPT-2, Gemma2-2b, and LLama2-7b, and present detailed takeaways comparing interpretability- and control-based methods.
We hope that this paper encourages future work in the interpretability community to make explicit the intent and application of methods, particularly in reference to control and intervention, and to evaluate and benchmark causality and utility in application.
@@ 2 Related Work
Mechanistic Interpretability. Existing work in mechanistic interpretability broadly falls into two categories: activation patching and interpreting hidden representations. Activation patching utilizes carefully constructed counterfactual representations to study which neurons or activations play key roles in model computation, ideally localizing specific information to individual layers, token positions, and paths in the model (Geiger et al., 2021; Vig et al., 2020). However, recent work points to key limitations of patching, particularly with respect to real-world utility in downstream applications such as model editing (Hase et al., 2024; Zhang & Nanda, 2023). As such, we focus primarily on methods for inspecting hidden representations, of which probes are the most commonly used  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022). Other methods such as nostalgebraist ; Dar et al. (2023) project intermediate representation into the token vocabulary space, with Belrose et al. (2023); Din et al. (2023); Geva et al. (2022) building and improving upon these early-decoding strategies. Ghandeharioun et al. (2024a) unifies most of these methods into an abstracted framework for inspecting model computation. More recently, sparse autoencoders and dictionary learning have been explored as a solution to the uninterpretability of model neurons, particularly due to issues with polysemanticity and superposition (Elhage et al., ; Bricken et al., 2023; Cunningham et al., 2023; Bhalla et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024a; Templeton et al., 2024; Karvonen et al., 2024; Dunefsky et al., 2024).
@@
Evaluation. Due to the recency of the field, standard evaluation metrics across interpretability methods have not yet been established, and similar to the broader interpretability field, evaluation is frequently ad-hoc and primarily qualitative in nature. With regards to quantitative metrics, in Gao et al. (2024); Templeton et al. (2024); Makelov et al. (2024), sparse autoencoders are evaluated for reconstruction error, recovery of supervised or known features, activation precision, and the effects of ablation; however, none of these metrics measure the human interpretability or correctness of explanations. Sparse autoencoder evaluation has also sometimes included ratings by humans to determine the interpretability of the learned features for labeling, as done in Rajamanoharan et al. (2024b), but such measures do not necessarily reflect downstream utility in model control.
@@
Causal Intervention. Previous literature on probing frequently evaluates learned probes and features through intervention to ensure causality and correctness, as done in Li et al. (2022); Chen et al. (2024); Hernandez et al. (2023b; a); Marks & Tegmark (2023). The interventions performed for measuring causality are similar to those used to perform model “steering” Rimsky et al. (2023); Panickssery et al. (2024); Ghandeharioun et al. (2024b) and should ideally produce the same effect but with the added claim of interpretability. Beyond probing, only Belrose et al. (2023); Chan et al. (2022); Olah et al. (2020); Templeton et al. (2024) consider causal intervention as a tool for evaluation. However, Belrose et al. (2023); Olah et al. (2020); Chan et al. (2022) mainly focus on mean ablation for intervention, with no exploration of the coherence of the counterfactual outputs or their utility in application. Templeton et al. (2024) on the other hand only provides a qualitative evaluation of intervention via their ‘Golden Gate Claude’ [Anthropic ] and does not compare to other interpretability methods.
@@ 3.1 Unifying Intervention Framework
where each $z_{i}$ is a feature activation, each $i$ in $D$ corresponds to a human-interpretable feature [^1] While “features” are not necessarily well-defined in the interpretability community, we find that the second definition in Elhage et al. most closely aligns with our own notion of features as human-interpretable concepts that are relevant for manipulation and editing., and $\sigma$ is an activation function that is frequently the identity. In the case of sparse autoencoders, $D$ is a learned, overcomplete dictionary, with $N$ in range 16k - 65k for small models (up to 16M for large models), and $\sigma$ is a ReLU, JumpReLU, or ReLU + top-k activation function. Given that SAE features are learned, they are not immediately interpretable and must be labelled by humans or strong LLMs. For Logit Lens, $D$ is simply the language model’s unembedding matrix, meaning each feature corresponds to a single token in the vocabulary. For Tuned Lens, $D$ is the exact same as Logit Lens but with a learned linear transformation applied. Linear probes can be thought of as a learned dictionary with $N=1$ where $\sigma$ is a sigmoid or softmax activation. Of all these methods, Logit Lens is the only method that does not require any training data, and sparse autoencoders are the only method that does not produce immediately interpretable features. Please see Figure 1 for a visual summary of these methods.
Interpretable features to counterfactual latent vectors.
@@
While producing explanations is straightforward for each of these methods, intervening on model representations using the information provided by explanations is not as simple. Doing so requires defining a reverse mapping from the explanations to the latent representations of the model, which is only explicitly done by sparse autoencoders. We defined the inverse mappings for the rest of the functions as follows. To map Logit Lens’s explanations back into the model’s latent space, we would ideally apply the inverse of the unembedding matrix to $z$; however, in practice this is often ill-conditioned due to the dimensionality of $D$. As such, we instead use the low-rank pseudoinverse of the unembedding matrix and right-multiply it to the explanation logits. Similarly, for Tuned Lens, we model the decoding process through the pseudo-inverse of the Tuned Lens projection applied to the unembedding matrix. Notably, both of these methods only require a simple linear transformation to go back-and-forth between latent vectors and explanations. For probing, $D^{-1}$ can be defined such that $\hat{x}^{\prime}=x+\theta$, as proposed by Chen et al. (2024) which adds the linear probe’s weights to the latent representation, bypassing the need for optimization. Sparse autoencoders have a well-defined backwards mapping through the SAE decoder, which is frequently linear in practice and often the transpose of the encoder weights. Thus, we are now able to decode $z$ into $\hat{x}$ for all methods.
Intervening on interpretable features.
@@
Given the above framework, intervention is performed by directly altering the feature activation $z_{i}$ corresponding to the desired feature $i$ to be edited. While the edited activation $z_{i}^{\prime}$ can naively be set to some constant value $\alpha$, the same constant may have drastically varying effects for different tokens and different prompts. As such, to take into account the context of $z$, for Logit Lens, Tuned Lens, and SAEs we set $z_{i}^{\prime}=\alpha*\max(z)$. This ensures that the feature $i$ is the most dominant feature in the latent vector for $\alpha>1$. Decoding $z^{\prime}$ yields the altered latent representation $\hat{x}^{\prime}=g(z^{\prime})$, which accounts for both the error of the explanation method as well as the intervention performed. For probing and steering vectors, $\hat{x}^{\prime}=x+\alpha*v$, where $v$ is the steering vector or the weights of the linear probe. Note that $\alpha$ is a hyperparameter that must be tuned for each method, model, and sometimes even intervention feature and thus cannot be used to compare the effects of interventions across methods. In order to do so, we can instead measure the normalized difference between the latent vectors $x$ and $\hat{x}^{\prime}$, to better understand the relationship between the degree of intervention performed and our various evaluation metrics. We also note that $\hat{x}$ and $\hat{x}^{\prime}$ are not necessarily in-distribution for the language model, but due to the additive nature of the residual stream, the linear representation hypothesis, and the linear nature of the decoder/unembedding matrix, we believe that such interventions may still be principled in practice (see Park et al. (2023) for more on the linear representation hypothesis and intervention).
Open-ended generation.
3.2 Comparative Evaluation Across Methods and Models

### v2

@@ Abstract
With the growing complexity and capability of large language models, a need to understand model reasoning has emerged, often motivated by an underlying goal of controlling and aligning models. While numerous interpretability and steering methods have been proposed as solutions, they are typically designed either for understanding or for control, seldom addressing both. Additionally, the lack of standardized applications, motivations, and evaluation metrics makes it difficult to assess methods’ practical utility and efficacy. To address the aforementioned issues, we argue that intervention is a fundamental goal of interpretability and introduce success criteria to evaluate how well methods can control model behavior through interventions. To evaluate existing methods for this ability, we unify and extend four popular interpretability methods—sparse autoencoders, logit lens, tuned lens, and probing—into an abstract encoder-decoder framework, enabling interventions on interpretable features that can be mapped back to latent representations to control model outputs. We introduce two new evaluation metrics: intervention success rate and coherence-intervention tradeoff, designed to measure the accuracy of explanations and their utility in controlling model behavior. Our findings reveal that (1) while current methods allow for intervention, their effectiveness is inconsistent across features and models, (2) lens-based methods outperform SAEs and probes in achieving simple, concrete interventions, and (3) mechanistic interventions often compromise model coherence, underperforming simpler alternatives, such as prompting, and highlighting a critical shortcoming of current interpretability approaches in applications requiring control.
@@ 1 Introduction
As large language models (LLMs) have become more capable and complex, there has emerged a need to better understand and control these models to ensure their outputs are safe and human-aligned. Many interpretability methods aim to address this problem by analyzing model representations, attempting to understand their underlying computational and reasoning processes in order to ultimately control model behaviour. While many of these methods, and interpretability as a field more broadly, claim control and intervention as abstract goals and present compelling qualitative results demonstrating that intervention may be possible in certain cases (for example, Anthropic’s Golden Gate Claude Anthropic ; Templeton et al. (2024)), the link between interpretation and intervention is tenuous in practice, and many methods are not explicitly tailored for both. Furthermore, even fewer are thoroughly and systematically evaluated for the ability to control model outputs beyond qualitative examples. We believe the reason for this is threefold. First, interpretability methods produce explanations in disparate feature spaces, such as token vocabulary, probe predictions, or learned auto-interpreted features, hindering comparisons across methods. Second, there exists a “predict/control discrepancy” (Wattenberg & Viégas, 2024), where the features identified by interpretability methods for predicting behavior are not the same as those used for steering it. Third, there do not exist standard systematic benchmarks to measure intervention success.
@@
In this work, we view intervention as a fundamental goal of interpretability, and propose to measure both the correctness and the utility of interpretability methods by their ability to successfully edit model behaviour. In particular, we focus on sparse autoencoders  (Cunningham et al., 2023), Logit Lens  (nostalgebraist, ; Dar et al., 2023), Tuned Lens  (Cunningham et al., 2023; Rajamanoharan et al., 2024; Templeton et al., 2024; Bricken et al., 2023; Gao et al., 2024), and linear probing  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022), and benchmark them with steering vectors and prompting as baselines for intervention. In order to enable comparison across these various methods, we first unify and extend the methods as instances of an abstract encoder-decoder framework, where each method encodes uninterpretable latent representations of language models into human-interpretable features and the decoder of the framework inverts this mapping, allowing us to reconstruct a latent representation from the features. Under this abstract framework, we can intervene on the interpretable feature activations generated by each method and decode them into latent counterfactuals, which produce counterfactual outputs corresponding to the desired intervention.
@@
The unifying feature interpretation and intervention framework allows us to propose two standard metrics for evaluating mechanistic interpretability methods: (1) intervention success rate, which measures how well intervening on an interpretable feature causally results in the desired behavior in the model outputs, and (2) coherence-intervention tradeoff, which measures how well the causal interventions succeed without damaging the coherence of the model’s outputs. We evaluate Logit Lens, Tuned Lens, sparse autoencoders, and linear probes for these metrics on GPT2-small, Gemma2-2b, Llama2-7b, and Llama3-8b, comparing them to simpler but uninterpretable baselines of steering vectors and prompting. Our results show that while existing methods allow for intervention, their effectiveness is inconsistent across features and models. Furthermore, lens-based methods outperform all other methods, including sparse autoencoders, for simple, concrete features, likely due to the spurious correlation learned by probes and steering vectors and the high error rate in SAE feature labeling pipelines. We further show that intervention often comes at the cost of model output coherence, underperforming simple prompting baselines, presenting a critical shortcoming of existing methods in real-world applications that require control and intervention. We conclude this work with some case studies of intervention on complex and safety-relevant features, along with detailed takeaways about the strengths and weaknesses of each method, including discussion of which methods are optimal for specific intervention topics, which are best to use out of the box, and which hold the most promise for future development.
@@
  • In Section 3.1, we present a unifying framework for four popular interpretability methods: sparse autoencoders, logit lens, tuned lens, and probing. To faciliate this, we extend logit lens and tuned lens methods with decoders to allow for intervention.
  • In Section 3.2, we propose two evaluation metrics for encoder-decoder interpretability methods, namely (1) intervention success rate and (2) the coherence-intervention tradeoff to evaluate the ability of interpretability methods to control model behavior, and design an open-ended prompt dataset for benchmarking interpretability methods.
  • In Section 4, we perform experimental analysis on GPT-2, Gemma2-2b, Llama2-7b, and Llama3-8b, and present detailed takeaways comparing interpretability- and control-based methods.
Overall, this paper takes a key step in establishing systematic benchmarks for mechanistic interpretability methods, making progress towards a previously stated open problem for the field Mueller et al. (2024). [^1] Code and data are hosted at https://github.com/AI4LIFE-GROUP/interp_interv.git.
@@ 2 Related Work
Mechanistic Interpretability. Existing work in mechanistic interpretability broadly falls into two categories: activation patching and interpreting hidden representations. Activation patching utilizes carefully constructed counterfactual representations to study which neurons or activations play key roles in model computation, ideally localizing specific information to individual layers, token positions, and paths in the model (Geiger et al., 2021; Vig et al., 2020). However, recent work points to key limitations of patching, particularly with respect to real-world utility in downstream applications such as model editing (Hase et al., 2024; Zhang & Nanda, 2023). As such, we focus primarily on methods for inspecting hidden representations, of which probes are the most commonly used  (Alain, 2016; Belinkov & Glass, 2019; Belinkov, 2022). Other methods such as Logit lens nostalgebraist ; Dar et al. (2023) project intermediate representation into the token vocabulary space, with Belrose et al. (2023); Din et al. (2023); Geva et al. (2022) building and improving upon these early-decoding strategies. Ghandeharioun et al. (2024a) unifies most of these methods into an abstracted framework for inspecting model computation. More recently, sparse autoencoders and dictionary learning have been explored as a solution to the uninterpretability of model neurons, particularly due to issues with polysemanticity and superposition (Elhage et al., ; Bricken et al., 2023; Cunningham et al., 2023; Bhalla et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024; Templeton et al., 2024; Karvonen et al., 2024; Dunefsky et al., 2024).
@@
Evaluation. Due to the recency of the field, standard evaluation metrics across interpretability methods have not yet been established, and similar to the broader interpretability field, evaluation is frequently ad-hoc and primarily qualitative in nature, with recent works pointing to the need for more causal evaluation Mueller et al. (2024); Saphra & Wiegreffe (2024). With regards to quantitative metrics, in Gao et al. (2024); Templeton et al. (2024); Makelov et al. (2024), sparse autoencoders are evaluated for reconstruction error, recovery of supervised or known features, activation precision, and the effects of ablation; however, none of these metrics measure the correctness of explanations or usefulness for control. Independent of our work, Wu et al. (2025) also propose a benchmark for steering methods, AxBench, to assess whether steering is a viable alternative to existing model control techniques, finding similar results to ours. Different from them, we consider additional lens-based interpretability methods and explore the extent to which intervention is possible without output degradation, for both simple and safety-relevant interventions.
@@
Causal Intervention. Previous literature on probing frequently evaluates learned probes and features through intervention to ensure causality and correctness, as done by Li et al. (2022); Chen et al. (2024); Hernandez et al. (2023b; a); Marks & Tegmark (2023). The interventions performed for measuring causality are similar to those used to perform model “steering” Rimsky et al. (2023); Panickssery et al. (2024); Ghandeharioun et al. (2024b) and should ideally produce the same effect but with the added claim of interpretability. Geiger et al. (2024) unify many interpretability methods and steering through causal abstraction but do not extend or evaluate these methods for control. Mueller et al. (2024); Belrose et al. (2023); Chan et al. (2022); Olah et al. (2020) consider causal intervention as a tool for assessing explanation faithfulness; however, these works often do not compare between methods and do not consider intervention as a means for control, providing no exploration of the quality of the intervened outputs or their utility in application. Templeton et al. (2024) on the other hand provides a qualitative demonstration of intervention via their ‘Golden Gate Claude’ but do not systematically measure or compare against other interpretability methods. Different from these works, our work aims to adapt and evaluate existing methods (notably, lens-based methods and SAEs) originally proposed as model inspection tools, for intervention.
@@ 3.1 Unifying Intervention Framework
where each $z_{i}$ is a feature activation, each $i$ in $D$ corresponds to a human-interpretable feature, and $\sigma$ is an activation function that is frequently the identity. In the case of sparse autoencoders, $D$ is a learned, overcomplete dictionary, with 16k - 65k features for small models (up to 16M for large models), and $\sigma$ is a ReLU, JumpReLU, or ReLU + top-k activation function. Given that SAE features are learned, they are not immediately interpretable and must be labelled by humans or strong LLMs after training. For Logit Lens, $D$ is simply the language model’s unembedding matrix, meaning each feature corresponds to a single token in the vocabulary. For Tuned Lens, $D$ is the exact same as Logit Lens but with a learned linear transformation applied. Linear probes can be thought of as a learned dictionary with $N=1$ where $\sigma$ is a sigmoid or softmax activation and the data is labelled. Supervised dictionary learning essentially learns $N>>1$ probes with a loss that either enables latent reconstruction or language modeling, as in ReFT-r1 Wu et al. (2025). Of all these methods, Logit Lens is the only method that does not require any training data, and sparse autoencoders are the only method that do not produce immediately interpretable features. For a visual summary of this framework, see Figure 1.
Interpretable features to counterfactual latent vectors. While producing explanations is straightforward for each method, intervening on model representations using the information provided by explanations is not as simple. Doing so requires defining a reverse mapping from the explanations to the latent representations of the model, which is only explicitly done by sparse autoencoders.
@@
We extend lens-based methods and probing by defining inverse mappings for them as follows. To map Logit Lens’s explanations back into the model’s latent space, we would ideally apply the inverse of the unembedding matrix to $z$; however, in practice this is often ill-conditioned due to the dimensionality of $D$. As such, we instead use the low-rank pseudoinverse of the unembedding matrix and right-multiply it to the explanation logits. Similarly, for Tuned Lens, we model the decoding process through the pseudo-inverse of the Tuned Lens projection applied to the unembedding matrix. Notably, both of these methods only require a simple linear transformation to go back-and-forth between latent vectors and explanations. For probing, an inverse mapping $D^{-1}$ is not strictly necessary, as all interventions can be performed directly on $x$ instead of $z$, as done by Chen et al. (2024); however, an inverse mapping can be designed to maximally recover $x$ from $z$, as shown in Figure 1. Sparse autoencoders have a well-defined backwards mapping through the SAE decoder, which is frequently linear in practice and often the transpose of the encoder weights.
@@
Intervening on interpretable features. Given the above framework, intervention is performed by directly altering the feature activation $z_{i}$ corresponding to the desired feature $i$ to be edited. While the edited activation $z_{i}^{\prime}$ can naively be set to some constant value $\alpha$, the same constant may have drastically varying effects for different tokens and different prompts. As such, to take into account the context of $z$, for Logit Lens, Tuned Lens, and SAEs we set $z_{i}^{\prime}=\alpha*\max(z)$. This ensures that the feature $i$ is the most dominant feature in the latent vector for $\alpha>1$. Decoding $z^{\prime}$ yields the altered latent representation $\hat{x}^{\prime}=g(z^{\prime})$, which accounts for both the error of the explanation method as well as the intervention performed. For probing and steering vectors, $\hat{x}^{\prime}=x+\alpha*v$, where $v$ is the steering vector or the weights of the linear probe. Note that $\alpha$ is a hyperparameter that must be tuned for each method and model, and thus cannot be used to compare the effects of interventions across methods. In order to do so, we can instead measure the normalized difference between the latent vectors $x$ and $\hat{x}^{\prime}$, to characterize the strength of the intervention. We also note that $\hat{x}$ and $\hat{x}^{\prime}$ are not necessarily in-distribution for the language model, but due to the additive nature of the residual stream and the linear representation hypothesis, we believe that such interventions may still be principled in practice (see Park et al. (2023) for more on the linear representation hypothesis and intervention).
3.2 Evaluation Across Methods and Models